A stone is thrown with  an initial speed of 4.9m/s from a bridge in vertically upward direction .It falls in ground in 2 sec.find height of bridge. Plz solve and eplain why answer is taken as positive if Accelaration due to gravity is taken negative 
 

Grade:11

3 Answers

Venkat
5 years ago
Take upward direction as positive.
s=ut+0.5at^2 \\ s= ut + 0.5(-a)t^2 \\ s=ut-0.5at^2 \\ \\h=4.9*2-0.5*9.8*4 = -9.8 m
As we have taken downward direction as positive, the negative answer say that height of building is measured upwards. Note that the distance cannot be negative as it is always greater than xero. Hence, the magnitude of height of building is 9.8
Khimraj
5 years ago
s=ut+0.5at^2 \\ s= ut + 0.5(-a)t^2 \\ s=ut-0.5at^2 \\ \\h=4.9*2-0.5*9.8*4 = -9.8 m

Venkat
5 years ago
@khimraj, dont partially copy my answers...do your  own and post it...i take time to solve and post it. So appreciate others work and be a gentleman...